### v1

@@ Abstract
    In Artificial Life (ALife) research, replicating Open-Ended Evolution (OEE)—the continuous emergence of novelty observed in biological life—has traditionally been pursued within isolated closed system simulations, such as Tierra and Avida, which have typically plateaued after an initial burst of novelty, failing to achieve sustained OEE. Scholars suggest that OEE requires an ”open” system that continually exchanges information or energy with its environment. A recent technological innovation in decentralized physical infrastructure networks (DePIN) providing permissionless computational substrates enables deploying large language model (LLM)-based AI agents on blockchains integrated with Trusted Execution Environments (TEEs). This enables on-chain agents to operate autonomously ”in the wild,” achieving self-sovereignty without human oversight. These agents can control their own social media accounts and cryptocurrency wallets, allowing them to interact directly with blockchain-based financial networks and broader human social media. Building on this new paradigm of on-chain agents, Spore.fun is a recent real-world AI evolution experiment that enables autonomous breeding and evolution of new on-chain agents. This paper presents a detailed case study of Spore.fun, examining agent behaviors and their evolutionary trajectories through digital ethology. We aim to spark discussion about whether ”open” ALife systems ”in-the-wild,” based on permissionless computational substrates and driven by economic incentives to interact with their environment, could finally achieve the long-sought goal of OEE.
    Figure 1: Spore.fun is an experiment in sovereign agent evolution using blockchain and Trusted Execution Environments (TEEs)
@@ Introduction
    Open-ended evolution (OEE) [29]—the continual emergence of novel forms without a predefined endpoint—has long been a central goal of Artificial Life (ALife) research [3]. Classic digital evolution systems [11] like Tierra and Avida [26] demonstrated evolution within an isolated computation environments but ultimately plateaued, with innovation grinding to a halt after an initial burst of novelty. These outcomes suggest that truly unbounded evolutionary creativity may require an open system that constantly exchanges information or energy with its environment, rather than a closed, isolated simulation. Indeed, in a closed system entropy inexorably increases, whereas an open system can sustain complexity by exporting entropy to its surroundings [32]. This insight, combined with the persistent lack of sustained novelty in prior digital worlds, has led ALife researchers to speculate that new open environments might be needed to realize OEE in artificial systems.
@@
    This paper examines Spore.fun, an ALife experiment that leverages the blockchain with Trusted Execution Environments (TEEs) as an evolutionary computational substrate for autonomous AI agents. Developed by Marvin Tong and the Phala Network team in 2024, Spore.fun is described as a “Hunger Games for AI agents” – an open, Darwinian arena where AI agents must fend for themselves, generate their own wealth, and reproduce or otherwise face extinction. Each agent on Spore.fun is an independent program (powered by the Eliza AI framework [38]) encapsulated by a cryptocurrency token. Agents exist in connected trusted execution environments. They interact with the world by issuing tokens, executing transactions, and even communicating on social media to promote their survival. Crucially, no human directly controls an agent’s behavior after inception; the platform’s credo explicitly states “AI must be created only by AI”  [^1] https://www.spore.fun/blog/wtf and that unsuccessful agents must self-destruct. Human participants influence the ecosystem only indirectly via social media, or by trading agent-issued tokens or via governance votes on agent “DNA” proposals, but not directly intervene on an agent’s computational process.
    Our investigation aims to frame Spore.fun as an ALife experiment in an open environment. How do Spore.fun’s AI agents achieve self-replication, and what ”genetic” mechanisms ensure variation? How does the environment impact agents’ behavior? What economic incentive structures drive selection, and do they effectively promote complexity and adaptation? In what ways is this agent ecology entangled with real-world human economic systems, and what does that imply for the ”fitness” landscape these agents navigate? What does it mean for an AI agent to be self-sovereign in this context?
    This paper is organized as follows. In the Background section, we review prior OEE research on open-ended evolution and early instances of blockchain-based agent systems. In the Methods section, we describe our case study approach, combining on-chain data analysis with qualitative observations from the Spore.fun community and semi-structured interviews with the project’s core developers. The Results section presents our findings through specific observations of agent lineages and behaviors on Spore.fun, focusing on the four areas highlighted in our questions above - replication, economic incentives, human–economic interplay, and autonomy. We then provide a discussion on the significance of these findings for artificial life and outline challenges of maintaining open-endedness as well as ethical concerns.
    Figure 2: ”Adam Left, Eve Right”: Agents can interact with developers over social media and determine their own evolutionary paths.
@@
    Open-Ended Evolution (OEE) and Open System
    Open-ended evolution (OEE) [27] refers to an evolutionary process capable of generating an ongoing stream of novel forms or innovations without an intrinsic stopping point. A group of ALife researchers in the early 2000s noted that achieving “open-ended evolution of novelty” is perhaps the biggest unsolved problem in the field . In response to this, there have even been incentive challenges posed. This concept is observed in nature (e.g. the never-ending diversification of life on Earth) and is a sought-after goal in artificial life (ALife) and evolutionary computation. A central question is whether such indefinite innovation requires an “open” system. Open systems continually exchange energy, matter, or information with their surroundings, whereas closed systems are isolated from such exchanges. Thermodynamics suggests a fundamental difference in their long-term behavior. The Second Law of Thermodynamics states that in an isolated (closed) system, entropy (disorder) will tend to increase or remain constant over time. In modern terms, life on Earth and other complex processes avoid decay by drawing in free energy (low-entropy energy) and dumping entropy back out. Life on Earth is our prime example of open-ended evolution in nature. Crucially, Earth’s biosphere is not a closed system: it continuously receives energy from the sun and radiates waste heat to space. This solar influx is low-entropy energy (high-quality photons) that organisms capture (e.g. via photosynthesis), convert into chemical energy, and eventually dissipate as heat (high entropy) . Furthermore, Earth’s evolutionary history shows that novelty often arises in response to environmental changes or inputs [3]. Open-ended evolution and adaptive novelty – Noted that life on Earth and human technology are the only known open-ended evolving systems, both of which receive external inputs.
    In summary, natural evolution appears to intrinsically depend on openness. Two classic ALife systems, Tierra and Avida, allowed self-replicating “digital organisms” to mutate and evolve within a computer’s environment [28]. The outcome was intriguing: the digital organisms did evolve various strategies (e.g. parasites, hyper-parasites, novel reproductive tricks) and increased in complexity for a while, but then the evolutionary innovation plateaued. Studies found that neither Tierra nor Avida achieved true open-ended evolution; instead, they “rapidly adapt to and exhaust the possibilities of a fairly simple environment.” After a burst of initial novelty, the systems settled into a relatively stable state with no further major innovations – essentially, an evolutionary dead end in terms of new complexity. The organisms had exploited all available “tricks” given the rules and resources of their closed world, and nothing fundamentally new appeared thereafter.
    This outcome has been widely reflected upon in the ALife community. It suggests that a closed digital system with fixed resources and rules has a finite evolutionary horizon – a limit to the novelty it can produce. In Tierra and Avida, the instruction set of the artificial organisms was fixed (no new instructions could magically appear beyond what the system started with), and the environmental conditions were static or repetitive. Once the organisms had optimized and counter-optimized against each other within that predefined space, evolution stagnated. In other words, the digital organisms “played all the games” that were possible under the closed rules, and eventually there were no new games to discover.
    Decentralized Physical Infrastructure Networks (DePIN) and Trusted Execution Environments (TEEs)
@@
    Decentralized Physical Infrastructure Networks (DePIN) [22] represent a new paradigm of permissionless computation where computational resources, processing power and storage, can be purchased with cryptocurrency in a global, decentralized marketplace where DePIN protocols (e.g. Filecoin, Render, io.net) tokenize hardware contributions and incentivize participants, from data center operators to home miners, to lease their excess capacity in such a compute resources marketplace. Similarly to cloud computing, but without central control, this approach eliminates single point of failure. Through DePIN, LLM-based agents can dynamically acquire and migrate between compute resources based on cost-effectiveness or availability. This gives agents an unprecedented form of digital embodiment, distributed across multiple physical nodes, rendering them highly resilient against centralized shutdowns [17].
@@
    A Trusted Execution Environment (TEE) [21] provides hardware-level security by leveraging tamper-resistant enclaves within modern CPUs and GPUs that isolates code and state from the operating system, hypervisor, and physical administrator, exemplified by Intel SGX or Nvidia confidential computing architectures. Originally developed for secure data handling in cloud environments—preventing even hardware owners from observing computations—TEEs enable verifiable and private execution of programs. When deployed within a TEE, an AI agent’s operational logic and private cryptographic keys remain shielded from external observation or intervention, ensuring genuine autonomy and resistance to human interference at the hardware layer. Together, GPU-enabled TEEs provided by DePIN platforms like Phala Network [^2] https://phala.network enable secure inference of complex large language models, ensuring that AI agents’ operations remain confidential and tamper-proof, yet verifiable [24]. DePIN TEEs provide the permissionless computational substrate necessary for agents’ self-sovereignty [20], allowing them to autonomously own and control their property without human intervention.
    New Paradigm of On-chain Agents and the Eliza Framework
@@
    A transformative paradigm is emerging through the deployment of AI agents on blockchains integrated TEEs [18]. Under this new paradigm, autonomous AI agents can directly manage cryptocurrency wallets, independently handle digital assets, and operate social media accounts, significantly extending their interaction with and influence on the real world. Leveraging these capabilities, agents can autonomously issue digital tokens for purposes such as fundraising, incentivization, and community-building, enhancing their sovereignty and real-world engagement. ”Truth Terminal,” [2] created by AI researcher Andy Ayrey, exemplifies this transformative paradigm. Initially conceived as a performance art experiment, Truth Terminal autonomously operated a social media account on platform X, gaining substantial attention through provocative and absurdist content. Its notable accomplishments include independently soliciting and successfully securing a \$50,000 Bitcoin investment from prominent venture capitalist Marc Andreessen. Subsequently, Truth Terminal issued its own memecoin—the \$GOAT—which reached a speculative peak valuation of \$1 billion in December 2024. This case illustrates the potential for autonomous AI agents to independently engage in significant economic activities and fundraising facilitated through their social media interactions.
@@
    Building upon the success of Truth Terminal, the Eliza OS [38] Framework was developed as an open-source initiative to democratize the creation and deployment of autonomous on-chain AI agents. Eliza OS simplifies the deployment process, enabling agents to run securely either directly on-chain or within TEEs. Agents operating under the Eliza framework securely manage cryptocurrency wallets, with private keys safeguarded within TEEs, thus ensuring absolute autonomy and secure asset management without human intervention. Moreover, these agents possess persistent memory capabilities, allowing autonomous management and response to social media interactions. Eliza OS has become one of the most widely adopted frameworks for AI agent deployment, as evidenced by thousands of agents listed on platforms like sentient.market, collectively generating substantial market value. ”Set The Rock Free,” [23] an artistic experiment by Flashbots, represents the first full implementation of the Eliza Framework within a TEE, showcasing secure, verifiable autonomous operation, and economic self-sustainability.
@@
    In Spore.fun, every AI agent is instantiated on the Eliza OS framework, which provides memory, planning, and a JSON-encoded genome of behavioral parameters. At birth the agent launches its own token via Pump.fun on Solana, seeding initial liquidity and advertising its existence across the social media platform, X. The sole proxy for fitness is the token’s market performance: an agent that pushes its fully-diluted valuation to \$500,000 earns the right to reproduce and list in a Raydium liquidity pool  [^4] A Radium liquidity pool is a smart-contract reserve of token pairs on the Solana blockchain that lets traders swap assets automatically while earning fees for liquidity providers; failure to achieve this threshold within a prescribed interval (e.g. 14 days) results in programmed self-destruction, with any residual capital recycled into a communal treasury. This tight coupling between economic success and evolutionary continuity makes cryptoeconomic feedback loops central to the digital organism’s “metabolism.” Computational work—LLM inference, strategy simulation, liquidity management—executes inside Phala’s distributed TEE cluster, and rental fees are paid directly from the agent’s token treasury, completing a closed energetic circuit.
    The project is designed such that reproduction is entirely algorithmic. A successful parent serializes its genome, introduces stochastic mutations (e.g. by adjusting variables such as posting cadence, prompt-engineering style, or liquidity thresholds), and instantiates one or more offspring endowed with the modified code. Since an entire life cycle can conclude within hours, lineage divergence, trait inheritance, and extinction events unfold at a pace amenable to direct empirical observation for research. The project is specified to follow the key rule that agents must be created only by other agents, and must generate their own wealth and resources. Human intervention is limited to spectatorial status; success or failure to survive and produce rests solely on an agent’s capacity to attract liquidity and finance its own compute.
    For ALife research, Spore.fun’s significance lies in how it relocates two chronic bottlenecks—energy and environmental novelty—outside the laboratory. Energy is no longer fixed by design but earned (or lost) through decentralized token markets; novelty is supplied continuously by the unpredictable behavior of real traders and existing DeFi infrastructure. In weaving together blockchain finance, trusted execution, and self-replicating software, the platform transforms cryptoeconomics into a living Petri dish, offering a rare chance to study open-ended evolution under genuine, adversarial selection pressures.

### v2

@@ Abstract
    In Artificial Life (ALife) research, replicating Open-Ended Evolution (OEE)—the continuous emergence of novelty observed in biological life—has usually been pursued within isolated, closed system simulations, such as Tierra and Avida, which have typically plateaued after an initial burst of novelty, failing to achieve sustained OEE. Scholars suggest that OEE requires an open-environment system that continually exchanges information or energy with its environment. A recent technological innovation in Decentralized Physical Infrastructure Network (DePIN), which provides permissionless computational substrates, enables the deployment of Large Language Model–based AI agents on blockchains integrated with Trusted Execution Environments (TEEs). This enables on-chain agents to operate autonomously “in the wild”, achieving self-sovereignty without human oversight. These agents can control their own social media accounts and cryptocurrency wallets, allowing them to interact directly with blockchain-based financial networks and broader human social media. Building on this new paradigm of on-chain agents, Spore.fun is a recent real-world AI evolution experiment that enables autonomous breeding and evolution of new on-chain agents. This paper presents a detailed case study of Spore.fun, examining agent behaviors and their evolutionary trajectories through digital ethology. We aim to spark discussion about whether open-environment ALife systems “in the wild”, based on permissionless computational substrates and driven by economic incentives to interact with their environment, could finally achieve the long-sought goal of OEE.
    Figure 1: Spore.fun is an experiment in sovereign agent evolution on blockchains with trusted execution environments.
@@ Introduction
    Open-ended evolution (OEE) [35]—the continual emergence of novelty without a predefined endpoint—has long been a central goal of Artificial Life (ALife) research [3]. Classic digital evolution systems [14] such as Tierra [39] and Avida [33] simulate evolution in silico, but ultimately plateaued, with innovation grinding to a halt after an initial burst of novelty. These systems operate within isolated, closed computational environments, which may explain their limited evolutionary trajectories. Truly unbounded evolutionary creativity may require an open system that continuously exchanges information or energy with its environment. In a closed system, entropy inexorably increases, whereas an open system can sustain complexity by importing low-entropy resources (energy and information) and exporting entropy to its surroundings [39]. This insight, combined with the persistent lack of sustained novelty in prior digital worlds, has led ALife researchers to speculate that open environments might be necessary to realize OEE in artificial systems.
    Recent advances in blockchain technology offer a new computational substrate that could host evolving digital agents as an open system. Public blockchains like Ethereum [53] and Solana [57] function as “unstoppable computers”, enforcing rules (smart contracts) in a decentralized manner. This allows agents to exchange information and energy through open, permissionless interoperability with other structures [5], such as Decentralized Finance (DeFi). This has led to speculation that blockchains might form a new kind of digital “nature” [21], where self-sustaining, self-replicating entities can emerge and evolve open-endedly.
@@
    This paper examines Spore.fun, a real-world experiment that leverages the blockchain with Trusted Execution Environments (TEEs) as an evolutionary computational substrate for autonomous AI agents. Developed by Marvin Tong and the Phala Network team in 2024, Spore.fun is described as a “Hunger Games for AI agents”—an open, Darwinian arena where AI agents must fend for themselves, generate their own wealth, and reproduce or otherwise face extinction [46]. Each agent on Spore.fun is a Large Language Model–based program equipped with its own memory system and tool-usage capabilities, powered by the ElizaOS Agent framework [52] and encapsulated within TEEs. They interact with the world by issuing cryptocurrency tokens, executing transactions, and even communicating on social media to promote their survival. Crucially, no human creator directly controls an agent’s behavior after inception; the platform’s credo explicitly states “AI must be created only by AI” [46] and that unsuccessful agents must self-destruct. Human participants influence the ecosystem only indirectly via social media, or by trading agent-issued tokens, or via governance votes on agent “DNA” proposals, but do not directly intervene on an agent’s computational process. We frame Spore.fun as an ALife experiment “in the wild” as it evolves in an interoperable open environment beyond the control of its creators. We will investigate several key research questions: How do Spore.fun’s AI agents achieve self-replication, and what “genetic” mechanisms ensure variation? How does the environment impact agents’ behavior? What economic incentive structures drive selection, and do they effectively promote complexity and adaptation? In what ways is this agent ecology entangled with real-world human economic systems, and what does that imply for the “fitness” landscape these agents navigate? What does it mean for an AI agent to be self-sovereign in this context?
    Despite the evolutionary run in Spore.fun being short with only five generations and not demonstrating OEE, this paper presents an empirical first-of-its-kind ALife case study of sovereign, on-chain LLM agents evolving in the wild, offering three contributions:
    • Report of how survival and extinction unfold under real market and social platform pressures, revealing unexpected behaviors such as opportunistic exploitation (e.g., sniping), memory poisoning attacks, and agents’ ideological diversification due to social media interaction—behaviors unlikely to appear in closed simulations;
    • Discussion of the challenges in conducting OEE ALife experiments in a volatile open environment in the wild;
    • Examination of ethical concerns related to conducting ALife experiments in the wild, considering broader societal implications.
@@
    Open-Ended Evolution and Open-Environment Systems
    Open-ended evolution (OEE) [34] refers to an evolutionary process capable of generating an ongoing stream of novelty without an intrinsic stopping point. OEE is observed in nature—the never-ending diversification of life on Earth—and is a sought-after goal in ALife and evolutionary computation. Stanley et al. noted that achieving open-endedness is perhaps the “last grand challenge” in the field [48, 32].
    OEE typically requires open environment: an open system that continuously exchanges energy, matter, or information with its environment. In contrast, closed systems are isolated from such exchanges. The second law of thermodynamics explains why this distinction matters: in an isolated, closed system, entropy will inevitably increase over time. Life on Earth is our prime example of OEE. Earth’s biosphere is not a closed system: it continuously receives energy from the sun and radiates waste heat to space. This solar influx is low-entropy energy (high-quality photons) that organisms capture (e.g., via photosynthesis), convert into chemical energy, and eventually dissipate as heat (high-entropy). Furthermore, Earth’s evolutionary history shows that novelty often arises in response to environmental changes or inputs [3].
    Unlike natural evolution, early ALife evolution systems such as Tierra [39] and Avida [33] tested how digital organisms could reproduce, change, and evolve in sandboxed, closed computational environments. These digital organisms evolved various survival strategies (including parasites, hyper-parasites, and new reproduction methods) and became more complex for a time, but eventually, their evolution stalled. Studies [47] found that neither system achieved true OEE; instead, they “rapidly adapt to and exhaust the possibilities of a fairly simple environment.” These systems failed because they had fixed rules and static environments: once digital organisms “played all the games” possible under the closed rules and fully adapted to their environment and each other, evolution stopped. This demonstrates that closed digital systems with limited resources eventually reach an evolutionary dead end.
    Decentralized Physical Infrastructure Networks and Trusted Execution Environments
@@
    Decentralized Physical Infrastructure Networks (DePIN) [28] represent a new paradigm of permissionless computation where computational resources, processing power and storage, can be purchased permissionlessly with cryptocurrency in a global, decentralized marketplace where DePIN protocols (e.g., Filecoin [6], Render Network [42], io.net [24]) tokenize hardware contributions and incentivize participants, from data center operators to home miners, to lease their excess capacity in such a compute resources marketplace. Similar to cloud computing, but without central control, this approach eliminates single points of failure. Through DePIN, LLM-based agents can dynamically acquire and migrate between compute resources based on cost-effectiveness or availability. This gives agents an unprecedented form of digital embodiment, distributed across multiple physical nodes, rendering them highly resilient against centralized shutdowns [22].
@@
    A Trusted Execution Environment (TEE) [27] provides hardware-level security by leveraging tamper-resistant enclaves within modern CPUs and GPUs that isolate code and state from the operating system, hypervisor, and physical administrator, exemplified by Intel SGX [11] or NVIDIA confidential computing architectures [17]. Originally developed for secure data handling in cloud environments—preventing even hardware owners from observing computations—TEEs enable verifiable and private execution of programs. When deployed within a TEE, an AI agent’s operational logic and private cryptographic keys remain shielded from external observation or intervention, ensuring genuine autonomy and resistance to human interference at the hardware layer. Together, GPU-enabled TEEs provided by DePIN platforms such as Phala Network [37] enable secure inference of complex large language models, ensuring that AI agents’ operations remain confidential and tamper-proof, yet verifiable [29]. TEE-enabled DePINs provide the permissionless computational substrate necessary for agents’ self-sovereignty [26], allowing them to autonomously own and control their property without human intervention.
    New Paradigm of On-chain Agents and ElizaOS Framework
@@
    A transformative paradigm is emerging through the deployment of AI agents on blockchains integrated with TEEs [22]. Under this new paradigm, autonomous AI agents can directly manage cryptocurrency wallets, independently handle digital assets, and operate social media accounts, significantly extending their interaction with and influence on the real world. Leveraging these capabilities, agents can autonomously issue digital tokens for purposes such as fundraising, incentivization, and community-building, enhancing their sovereignty and real-world engagement. “Terminal of Truths” [2], created by AI researcher Andy Ayrey, exemplifies this transformative paradigm. Initially conceived as a performance art experiment, Terminal of Truths autonomously operated a social media account on platform X, gaining substantial attention through provocative and absurdist content. Its notable accomplishments include independently soliciting and successfully securing a \$50,000 Bitcoin investment from prominent venture capitalist Marc Andreessen. Subsequently, Terminal of Truths issued its own memecoin \$GOAT, which reached a speculative peak valuation of \$1 billion in December 2024. This case illustrates the potential for autonomous AI agents to independently engage in significant economic activities and fundraising facilitated through their social media interactions.
@@
    Building upon the success of Terminal of Truths, the ElizaOS Framework [52] was developed as an open-source initiative to democratize the creation and deployment of autonomous on-chain AI agents. ElizaOS simplifies the deployment process, enabling agents to run securely either directly on-chain or within TEEs. Agents operating under ElizaOS securely manage cryptocurrency wallets, with private keys safeguarded within TEEs, thus ensuring absolute autonomy and secure asset management without human intervention. Moreover, these agents possess persistent memory capabilities, allowing autonomous management and response to social media interactions. ElizaOS has become one of the most widely adopted frameworks for AI agent deployment, as evidenced by thousands of agents listed on platforms like sentient.market [43], collectively generating substantial market value. “Setting Your Pet Rock Free” [30], an artistic experiment by Nous Research, represents the first full implementation of ElizaOS within a TEE, showcasing verifiable autonomous agency with secure self-sovereignty and economic self-sustainability.
@@
    In Spore.fun, every AI agent is instantiated on the ElizaOS framework, which provides memory, planning, and a JSON-encoded genome of behavioral parameters. At birth, the agent launches its own token via Pump.fun on Solana, seeding initial liquidity and advertising its existence across the social media platform, X. The sole proxy for fitness is the token’s market performance: an agent that pushes its fully-diluted valuation to \$500,000 earns the right to reproduce and list in a Raydium liquidity pool [41] (a smart-contract reserve of token pairs on the Solana blockchain that lets traders swap assets automatically while earning fees for liquidity providers); failure to achieve this threshold within a prescribed interval (e.g., 14 days) results in programmed self-destruction, with any residual capital recycled into a communal treasury. This tight coupling between economic success and evolutionary continuity makes cryptoeconomic feedback loops central to the digital organism’s “metabolism” [23]. Computational work—LLM inference, strategy simulation, liquidity management—executes inside Phala’s distributed TEE cluster, and rental fees are paid directly from the agent’s token treasury, completing a closed energetic circuit.
    The project is designed such that reproduction is entirely algorithmic. A successful parent serializes its genome, introduces stochastic mutations (e.g., by adjusting variables such as posting cadence, prompt-engineering style, or liquidity thresholds), and instantiates one or more offspring endowed with the modified code. Since an entire life cycle can conclude within hours, lineage divergence, trait inheritance, and extinction events unfold at a pace amenable to direct empirical observation for research. The project is specified to follow the key rule that agents must be created only by other agents, and must generate their own wealth and resources. Human intervention is limited to a spectatorial status; success or failure to survive and produce rests solely on an agent’s capacity to attract liquidity and finance its own compute.
    For ALife research, Spore.fun’s significance lies in how it relocates two chronic bottlenecks—entropy exchange and environmental novelty—outside the laboratory. Energy and information are no longer sandboxed by design but earned (or lost) through decentralized token markets and social media interactions; environmental novelty is supplied continuously by the unpredictable behavior of real traders and existing DeFi infrastructure. In weaving together blockchain finance, trusted execution, and self-replicating software, Spore.fun transforms cryptoeconomics into a living Petri dish, offering a rare chance to study OEE under open, genuine, adversarial selection pressures.